# Elastic scattering at $\sqrt{s}=6$ GeV up to $\sqrt{s}=13$ TeV

### v1

@@ VIII Proton-neutron elastic scattering
We take in our analysis $24$ sets of the proton-neutron experimental data from $\sqrt{s}>4.5$ GeV up to maximum energy $\sqrt{s}>27.5$ GeV where they correspond to experimental data with a total number of experimental points $N=526$. As in the case of $pp$ scattering, we include in our fitting procedure only statistical errors with taken into account the systematic errors as additional normalization of a separate set. We take the amplitudes of the model obtained in the case of $pp$ and $p\bar{p}$ scattering and fix the parameters of the main terms and two anomalous terms, except for the terms of the second Reggions. The electromagnetic amplitudes are removed in this case. As a result, in the $pn$ case, only parameters of second effective Reggion are obtained from the fitting procedure. The $\sum_{i,j}\chi^{2}_{i,j}=567$. The corresponding obtained description of the differential cross section are present in Figs. 18, 19. It is clear that the HEGS model very well works in the case of proton-neutron elastic scattering. As in the case of $pn$ scattering, there are many experimental data at small momentum transfer, where the contributions of our two anomalous terms are important, the successful description of the differential cross section confirms the existence of these terms.
Figure 20: The analyzing power $A_{N}$ of pp - scattering calculated: a) at $\sqrt{s}=3.62$GeV (small $t$) (the experimental data [102]), and b) at $\sqrt{s}=3.62$GeV (larger $t$).
Figure 21: The analyzing power $A_{N}$ of pp - scattering calculated: a) at $\sqrt{s}=9.2$GeV, (the experimental data [102]), and b) at $\sqrt{s}=13.7$GeV (points - the experimental data [102]).
Figure 22: The analyzing power $A_{N}$ of pp - scattering calculated: a) at $\sqrt{s}=9.2$GeV, (the experimental data [102]), and b) at $\sqrt{s}=13.7$GeV (points - the experimental data [102]).
Figure 23: The analyzing power $A_{N}$ of pp - scattering calculated: a) at $\sqrt{s}=19.4$GeV (the experimental data [102]), and b) at $\sqrt{s}=23.4$GeV (points - the existing experimental data [102])
@@ IX Polarization effects in proton-proton elastic scattering
The corresponding spin-correlation values presented in eq.(30).
Neglecting the $\Phi_{2}(s,t)-\Phi_{4}(s,t)$ contribution, the spin correlation parameter $A_{N}(s,t)$ can be written taking into account the phases of separate spin non-flip and spin-flip amplitudes as $\varphi_{nf}(s,t),\varphi_{sf}(s,t)$ the analysis power is
@@
It is usually assumed that the imaginary and real parts of the spin-non-flip amplitude have the exponential behavior with the same slope, and the imaginary and real parts of the spin-flip amplitudes, without the kinematic factor $\sqrt{|t|}$ [100], are proportional to the corresponding parts of the non-flip amplitude. That is not so as regards the $t$ dependence shown in Ref. [70], where $F^{fl}_{h}$ is multiplied by the special function dependent on $t$. Moreover, one mostly takes the energy independence of the ratio of the spin-flip parts to the spin-non-flip parts of the scattering amplitude. All this is our theoretical uncertainty [87, 86].
In [88, 89] on the basis of generalization of the constituent-counting rules of the perturbative QCD, the proton current matrix elements $J^{\pm\delta\delta}_{p}$ for a full set of spin combinations corresponding to the number of the spin-flipped quarks were calculated. This leads to part of the spin-flip amplitude
@@
Of course, at lower energies we need to take into account the energy dependence parts of the spin-flip amplitudes. So the form of the spin-flip amplitude is determined as
@@
Our calculation for $A_{N}(t)$ is shown in Fig. 21 a,b at $\sqrt{s}=4.9$ GeV and $\sqrt{s}=6.8$ GeV. For our high energy model it is a very small energy. However, the description of the existing data is sufficiently good. At these energies, the diffraction minimum is practically overfull by the real part of the spin-non-flip amplitude and the contribution of the spin-flip-amplitude; however, the $t$-dependence of the analysing power is very well reproduced in this region of the momentum transfer. Note that the magnitude and the energy dependence of this parameter depend on the energy behavior of the zeros of the imaginary-part of the spin-flip amplitude and the real-part of the spin-nonflip amplitude. Figure 22 shows $A_{N}(t)$ at $\sqrt{s}=9.2$ GeV and $\sqrt{s}=13.7$ GeV. At these energies the diffraction minimum deepens and its form affects the form of $A_{N}(t)$. At last, $A_{N}(t)$ is shown at large energies $\sqrt{s}=19.4$ GeV and $\sqrt{s}=23.4$ GeV in Fig. 23. The diffraction dip in the differential cross section has a sharp form and it affects the sharp form of $A_{N}(t)$. The maximum negative values of $A_{N}$ coincide closely with the diffraction minimum.
We have found that the contribution of the spin-flip to the differential cross sections is much less than the contribution of the spin-nonflip amplitude in the examined region of momentum transfers from these figures; $A_{N}$ is determined in the domain of the diffraction dip by the ratio
@@
The size of the analyzing power changes from $-45\%$ to $-50\%$ at $\sqrt{s}=50$ GeV up to $-25\%$ at $\sqrt{s}=500$ GeV. These numbers give the magnitude of the ratio Eq.(53) that does not strongly depend on the phase between the spin-flip and spin-nonflip amplitudes. This picture implies that the diffraction minimum is mostly filled by the real-part of the spin-nonflip amplitude and that the imaginary-part of the spin-flip amplitude increases in this domain as well.
@@ X Conclusions
Practically for first time, a simultaneously research of proton-proton, proton-antiproton and proton-neutron elastic scattering has been carried out in a wide energy ((from $3.6$ GeV up to $13$ TeV) and momentum transfer region (from $|t|=2.10^{-4}$ GeV2 up to $|t|=14$ GeV2. In the fitting procedure we used only statistical errors. Systematic errors, which are mostly determined by indefinite of a luminosity, were taken into account as additional normalization coefficient. As a result, a wide range of possible forms of the scattering amplitudes are pretty decreases. As a result, a simultaneously description of the cross sections and spin correlation parameter of different nucleon-nucleon reactions, including 90 sets of experimental data, with the total number of data $N=4326$ gives very reasonably $\sum_{i,j}\chi^{2}_{i,j}=4826$. The $pn$ case with $526$ experimental data, where the basic parameters were fixed from $pp$ and $p\bar{p}$ scattering, $\sum_{i,j}\chi^{2}_{i,j}=585$.
@@
Our analysis is carried out by using a successful development of the HEGS model which can be applied in the wide energy and momentum transfer regions. The model of hadron interaction is based on the analyticity of the scattering amplitude with taking into account the hadron structure, which is represented by GPDs. Different origins of the non-linear behavior of the slope of the scattering amplitude are analyzed. The possible contribution of a meson threshold is compared with different forms of the approximations for a non-linear slope at small momentum transfer.
@@
The relative contributions of the possible different part of the scattering amplitude were especially analyzed. It is remarkable that in the model the main pomeron and odderon amplitudes have the same intercept. This leads after eikonalization to the $\ln^{2}(s)$ energy dependence. In this sense, we have some case of maximal Odderon. In the model, the odderon amplitude has a special kinematic factor and does not give a visible contribution at zero momentum transfer.
@@
It was found that the new anomalous term with a large slope has the complicated logarithmic energy dependence and has the cross even properties. Hence, it is part of the pomeron amplitude and is also proportional to charge distributions. Our analysis of the contribution of the so-called hard pomeron with a large intercept does not show a visible contribution of this term. The second additional term, which represents the additional oscillation properties of the scattering amplitude at small momentum transfer with cross-odd properties has an logarithmic energy dependence and is proportional to gravitomagnetic form factor. Hence, it belongs to the odderon contribution in the scattering amplitude.
@@
Our work supports that GPDs reflect the basic properties of the hadron structure and provide some bridge between many different reactions. The determined new form of the momentum transfer dependence of GPDs allows one to obtain different form factors, including Compton form factors, electromagnetic form factors, transition form factor, and gravitational form factors. The chosen form of the $t$-dependence of GPDs of the pion (the same as the $t$-dependence of the nucleon) allow us to describe the electromagnetic and GFF of pion and pion-nucleon scattering.
The impact of the different $t$ dependence of the real and imaginary parts of the elastic scattering amplitude at small $t$ should be noted. The dispersion relation shows that the real part has zero at small $t$ , approximately in the domain $-t=0.1$; of course this depends on the energy. Hence, the contrary behavior of the real part (growth at small $t$ so-called ”peripheral case” of the phase of the scattering amplitude) which is examined in [32] has no physical meaning.

### v3

@@ VIII Proton-neutron elastic scattering
We take in our analysis $24$ sets of the proton-neutron experimental data from $\sqrt{s}>4.5$ GeV up to maximum energy $\sqrt{s}>27.5$ GeV where they correspond to experimental data with a total number of experimental points $N=526$. As in the case of $pp$ scattering, we include only statistical errors in our fitting procedure taking into account the systematic errors as additional normalization of a separate set. We take the amplitudes of the model obtained in the case of $pp$ and $p\bar{p}$ scattering and fix the parameters of the main terms and two anomalous terms, except for the terms of the second Reggions. The electromagnetic amplitudes are removed in this case. As a result, in the $pn$ case, only the parameters of the second effective Reggion are obtained from the fitting procedure. Only in the Pomeron-neutron vertex $F_{Pom2}^{Born}n$ is taken in the standard dypole form with free $\Lambda_{pom-n}$. Its value is obtained to be slightly above the standard electromagnetic $\Lambda_{pom-n}^{2}=0.82$ GeV2. The obtained $\sum_{i,j}\chi^{2}_{i,j}=567$. The corresponding obtained description of the differential cross section is represent in Figs. 4, 21, 22. It is clear that the HEGS model works very well in the case of proton-neutron elastic scattering. In Fig. 22, the differential cross section of $pn$ elastic scattering is represented for a large momentum transfer at energies $\sqrt{s}=26.19$ GeV and $\sqrt{s}=19.4$ GeV. At $\sqrt{s}=26.19$ GeV (Fig. 22[left]) the diffractive dip-bump structure is clearly represented whereas at $\sqrt{s}=19.4$ GeV (Fig. 22[right]) this structure is being leveled out and practically disappears at more lower energies (see Fig. 23). As in the case of $pn$ scattering, there are many experimental data at small momentum transfer, where the contributions of our two anomalous terms are important; the successful description of the differential cross section confirms the existence of these terms.
Figure 24: The analyzing power $A_{N}$ of pp - scattering is calculated: [left] at $\sqrt{s}=3.62$ GeV; (small $t$), [right] (larger $t$).
Figure 25: The analyzing power $A_{N}$ of pp - scattering calculated: [left] at $\sqrt{s}=4.9$GeV, (the experimental data [61]), and [right] at $\sqrt{s}=6.8$GeV (points - the experimental data [61]).
Figure 26: The analyzing power $A_{N}$ of pp - scattering calculated: [left] at $\sqrt{s}=9.2$GeV (the experimental data [61]) and [right] at $\sqrt{s}=13.7$GeV (points - the experimental data [61]).
Figure 27: The analyzing power $A_{N}$ of pp - scattering calculated: [left] at $\sqrt{s}=19.4$GeV (the experimental data [61]),
and [right] at $\sqrt{s}=23.4$GeV (points - the existing experimental data [61])
@@ IX Polarization effects in proton-proton elastic scattering
The corresponding spin-correlation values are presented in eq.(29).
Neglecting the $\Phi_{2}(s,t)-\Phi_{4}(s,t)$ contribution, the spin correlation parameter $A_{N}(s,t)$ can be written taking into account the phases of separate spin non-flip and spin-flip amplitudes as $\varphi_{nf}(s,t),\varphi_{sf}(s,t)$ the analyzing power is
@@
It is usually assumed that the imaginary and real parts of the spin-non-flip amplitude have the exponential behavior with the same slope, and the imaginary and real parts of the spin-flip amplitudes, without the kinematic factor $\sqrt{|t|}$ [101], are proportional to the corresponding parts of the non-flip amplitude. That is not so as regards the $t$ dependence shown in Ref. [69], where $F^{fl}_{h}$ is multiplied by a special function dependent on $t$. Moreover, one mostly takes the energy independence of the ratio of the spin-flip parts to the spin-non-flip parts of the scattering amplitude. All this is our theoretical uncertainty [103, 102].
In [104, 105], on the basis of generalization of the constituent-counting rules of the perturbative QCD, the proton current matrix elements $J^{\pm\delta\delta}_{p}$ for a full set of spin combinations corresponding to the number of the spin-flipped quarks were calculated. This leads to part of the spin-flip amplitude
@@
Of course, at lower energies we need to take into account the energy-dependent parts of the spin-flip amplitudes. So the form of the spin-flip amplitude is determined as
@@
The minimal energy at which the analyzing power $A_{N}(s,t)$ of pp - scattering examined in the HEGS model is $\sqrt{s}=3.62$ GeV. For our high energy model it is a very small energy. These results are shown in Fig. 24 at small $t$ [left] and at larger $t$ (right). These model calculations are compared with the results of four experiments. Obviously, we obtain a very good description of the experimental data. At these energies, the diffraction minimum is practically overfull by the real part of the spin-non-flip amplitude and the contribution of the spin-flip-amplitude; however, the $t$-dependence of the analysing power is very well reproduced in this region of the momentum transfer.
Our calculation for $A_{N}(t)$ is shown in Fig. 25 (left, right) at $\sqrt{s}=4.9$ GeV and $\sqrt{s}=6.8$ GeV. The description of the existing data is sufficiently good. Note that the magnitude and the energy dependence of this parameter depend on the energy behavior of the zeros of the imaginary-part of the spin-flip amplitude and the real-part of the spin-nonflip amplitude. Figure 25 shows $A_{N}(t)$ at $\sqrt{s}=9.2$ GeV and $\sqrt{s}=13.7$ GeV. At these energies the diffraction minimum deepens and its form affects the form of $A_{N}(t)$. At last, $A_{N}(t)$ is shown at large energies $\sqrt{s}=19.4$ GeV and $\sqrt{s}=23.4$ GeV in Fig. 27. The diffraction dip in the differential cross section has a sharp form and it affects the sharp form of $A_{N}(t)$. The maximum negative values of $A_{N}$ coincide closely with the diffraction minimum.
We have found that the contribution of the spin-flip to the differential cross sections is much less than the contribution of the spin-nonflip amplitude in the examined region of momentum transfers. $A_{N}$ is determined in the domain of the diffraction dip by the ratio
@@
The size of the analyzing power changes from $-45\%$ to $-50\%$ at $\sqrt{s}=50$ GeV up to $-25\%$ at $\sqrt{s}=500$ GeV. These numbers give the magnitude of the ratio, eq.(51), that does not strongly depend on the phase between the spin-flip and spin-nonflip amplitudes. This picture implies that the diffraction minimum is mostly filled by the real-part of the spin-nonflip amplitude and that the imaginary-part of the spin-flip amplitude increases in this domain as well.
@@ X Conclusions
Practically for the first time, a simultaneous research of proton-proton, proton-antiproton and proton-neutron elastic scattering has been carried out in a wide energy (from $3.6$ GeV up to $13$ TeV) and momentum transfer region (from $|t|=2.10^{-4}$ GeV2 up to $|t|=14$ GeV2). In the fitting procedure we used only statistical errors. Systematic errors, which are mostly determined by indefiniteness of luminosity, take into account as additional normalization coefficient. As a result, a wide range of possible forms of the scattering amplitudes decrease considerably. In this paper a simultaneous description of the cross sections and spin correlation parameter of different nucleon-nucleon reactions, including 90 sets of experimental data, with the total number of data $N=4326$ gives a very reasonable $\sum_{i,j}\chi^{2}_{i,j}=4826$. The $pn$ case with $526$ experimental data, where the basic parameters were fixed from $pp$ and $p\bar{p}$ scattering, $\sum_{i,j}\chi^{2}_{i,j}=585$.
@@
Our analysis is carried out by using a successful development of the HEGS model which can be applied in a wide energy and momentum transfer regions. The model of hadron interaction is based on the analyticity of the scattering amplitude with taking into account the hadron structure, which is represented by GPDs. Different origins of the non-linear behavior of the slope of the scattering amplitude are analyzed. The possible contribution of a meson threshold is compared with different forms of the approximations for a non-linear slope at a small momentum transfer.
@@
The relative contributions of the possible different part of the scattering amplitude were especially analyzed. It is remarkable that in the model the main pomeron and odderon amplitudes have the same intercept. After eikonalization this leads to the $\ln^{2}(s)$ energy dependence. In this sense, we have some case of maximal Odderon. In the model, the odderon amplitude has a special kinematic factor and does not give a visible contribution at a zero momentum transfer.
@@
It was found that the new anomalous term with a large slope has the complicated logarithmic energy dependence and has the cross even properties. Hence, it is part of the pomeron amplitude and is also proportional to charge distributions. Our analysis of the contribution of the so-called hard pomeron with a large intercept does not show a visible contribution of this term. The second additional term, which represents the additional oscillation properties of the scattering amplitude at a small momentum transfer with the cross-odd properties, has a logarithmic energy dependence and is proportional to the gravitomagnetic form factor. Hence, it belongs to the odderon contribution to the scattering amplitude.
As was noted in many models, the oscillations of the scattering amplitude are connected with the broken Pomeranchuk theorem and the scattering amplitude grows to a maximal possible extent but not breaking the Froissart boundary. In our opinion that oscillation of the scattering amplitude is connected with the behavior of hadron interaction potential at large distances [31, 106, 107]. Note that in [108] a quark-antiquark potential as function of the separation distance $r$ between the particles for the perturbative case in the framework of the Refined Gribov-Zwanziger Model (RGZ) was calculated. It includes a combination of oscillating terms. This can compared to the Perturbative approach at short distances and Lattice QCD in the IR regime. So the researches into the properties of the oscillations term can give a very important information about the hadron interactions at large distances. Of course, this requires further theoretical and experimental works.
@@
Our work supports that GPDs reflect the basic properties of the hadron structure and provide a bridge between many different reactions. The determined new form of the momentum transfer dependence of GPDs allows one to obtain different form factors, including Compton form factors, electromagnetic form factors, transition form factor, and gravitational form factors. The chosen form of the $t$-dependence of GPDs of the pion (the same as the $t$-dependence of the nucleon) allows us to describe the electromagnetic and gravitomagnetic form factors of pion and pion-nucleon scattering.
The impact of the different $t$ dependence of the real and imaginary parts of the elastic scattering amplitude at small $t$ should be noted. The dispersion relation shows that the real part has zero at small $t$ , approximately in the domain $-t=0.1$; of course this depends on the energy. Hence, the contrary behavior of the real part (grows at small $t$, the so-called ”peripheral case” of the phase of the scattering amplitude) which is examined in [36] has no physical meaning.